\begin{thebibliography}{ABCD}


%\bibitem{CS}
%{\sc C, Alberto; S, Paolo:} Uniqueness of dg enhancements for  the derived category of a Grothendieck category. arXiv: 1507.05509v4. 2017.
\bibitem{Beilinson}
{\sc A. Beilinson}:
\emph{Remarks on Grothendieck's standard conjectures},  Regulators, 25–32, Contemp. Math., 571, Amer. Math. Soc., Providence, RI, 2012.

\bibitem{Cristina}
{\sc R. M. Cristina}:
\emph{A threefold of general type with {$q_1=q_2=p_g=P_2=0$}}. Acta Appl. Math. 2003, 133--150.

\bibitem{Friedlander1} E. Friedlander, {\sl Algebraic cycles, Chow
varieties, and Lawson homology.}  Compositio Math. 77 (1991), no. 1,
55--93.

\bibitem{Friedlander2} E. Friedlander,
\emph{Filtrations on algebraic cycles and homology. }
Ann. Sci. \'Ecole Norm. Sup. (4) 28 (1995), no. 3, 317--343.

%\bibitem{Friedlander-Haesemesyer-Walker} E. Friedlander;
%C. Haesemeyer and M. Walker, {\sl Techniques, computations, and
%conjectures for semi-topological K-theory.}
%Math. Ann. 330 (2004), no. 4, 759--807.


%%\bibitem{Friedlander-Gabber} E. Friedlander and  O. Gabber, {\sl Cycle
%spaces and intersection theory. Topological methods in modern
%mathematics} (Stony Brook, NY, 1991), 325--370, Publish or Perish,
%Houston, TX, 1993.

\bibitem{Friedlander-Mazur} E. Friedlander and B. Mazur,
{\sl Filtrations on the homology of algebraic varieties. With an
appendix by Daniel Quillen.}  Mem. Amer. Math. Soc. 110 (1994), no.
529


%\bibitem{FW}
%{\sc E. M. Friedlander and M. E. Walker}: Comparing K-theories for complex varieties, Amer. J. Math. 123, no. 5,
%779-810.  2001.


\bibitem{Hu}
{\sc W. Hu}: {\sl Birational invariants defined by Lawson homology}. Michigan Math. J. 60 (2011), 331-354.


%\bibitem{Hu2} {\sc W. Hu}:
%Lawson homology for abelian varieties.
%Forum Math.  27  (2015),  no. 2, 1249–1276.

%\bibitem{Hu-Li}
%{\sc W. Hu, L. Li}:
%Lawson homology, morphic cohomology and Chow motives.
%Math. Nachr. 284 (2011), no. 8-9, 1024-1047.


\bibitem{Kleiman}
{\sc S. Kleiman}: The standard conjectures. Motives (Seattle, WA, 1991), 3–20, Proc. Sympos. Pure Math., 55, Part 1, Amer. Math. Soc., Providence, RI, 1994.

%\bibitem{K}
%{\sc K. K\"{u}nnemann}: A Lefschetz decomposition for Chow motives of abelian schemes.



\bibitem{Lawson1}
{\sc H. B. Lawson, Jr.}: {\sl Algebraic cycles and homotopy theory.}, Ann.
of Math. {\bf 129}(1989), 253-291.

\bibitem{Lewis}
{\sc J.D. Lewis}: A survey of the Hodge conjecture. Universit\'{e} de Montr\'{e}al, Centre de Recherches Math\'{e}matiques, Montreal, QC. 1991.

\bibitem{Tankeev}
{\sc Tankeev, S. G.}: On the standard conjecture of Lefschetz type for complex projective threefolds. II. Izv. Math. 75 (2011),  no. 5, 1074--1062.

\end{thebibliography}

\documentclass[12pt]{amsart}
\usepackage[a4paper, left=3cm, right=3cm, height=22cm]{geometry}
\usepackage{amsmath,amsfonts,amsthm, amssymb}
\usepackage[arrow,matrix, all,cmtip]{xy}
\usepackage{enumerate}

\usepackage[usenames]{color}
\usepackage{amssymb}

\newtheorem{theorem}{Theorem}[section]
\numberwithin{equation}{section}

\theoremstyle{remark}

\newtheorem{conjecture}[equation]{Conjecture}

\def\Z{{\mathbb{Z}}}

\usepackage[pdfauthor={Jin Cao}, %
				pdftitle={Cdga}%
			dvips,colorlinks=true]{hyperref}
\hypersetup{
	bookmarksnumbered=true,
	linkcolor=black,
}
\newcounter{elno}

\begin{document}
\author{Jin Cao and Wenchuan Hu}
\title{The equivalence of Friedlander-Mazur  and Standard conjectures for threefolds}
\date{\today}

\address{
Jin Cao\\
Yau Mathematical Sciences Center,
Tsinghua University, Beijing, China
}
\email{caojin@mail.tsinghua.edu.cn}

\address{Wenchuan Hu,
School of Mathematics,
Sichuan University,
Chengdu,
China
}
\email{huwenchuan@gmail.com}


\begin{abstract}
 We show that the Friedlander-Mazur conjecture holds for a complex smooth projective variety $X$ of dimension three implies the standard conjectures hold for $X$. This together with a result of Friedlander yields the equivalence of the two conjectures in dimension three.
 From this we provide some new examples whose standard conjectures hold.
\end{abstract}
\maketitle
\pagestyle{myheadings}
 \markright{The equivalence of Friedlander-Mazur  and Standard conjectures for threefolds}

\section{The main result}
Let $X$ be a complex projective variety of dimension $n$. The Lawson homology
$L_pH_k(X)$ of $p$-cycles for a projective variety is defined by
$$L_pH_k(X) := \pi_{k-2p}({\mathcal Z}_p(X)) \quad for\quad k\geq 2p\geq 0,$$
where ${\mathcal Z}_p(X)$ is the space of algebraic $p$-cycles on $X$ provided with a natural topology (see
\cite{Friedlander1}, \cite{Lawson1}).

In \cite{Friedlander-Mazur}, Friedlander and Mazur define a cycle class map
\begin{equation}\label{eq01}
\Phi_{p,k}:L_pH_{k}(X)\rightarrow H_{k}(X,\Z)
\end{equation}
for all $k\geq 2p\geq 0$. Suslin conjectures  that $\Phi_{p,k}:L_pH_{k}(X)\rightarrow H_{k}(X,\Z)$ is an injection for $k=n+p-1$ and an isomorphism for $k\geq n+p$.

From now on, all the (co)homology have the rational coefficients. Let $X$ be a complex projective variety of dimension 3.
In this case, the second author showed in \cite{Hu} that the Friedlander-Mazur conjecture \cite[\S 7]{Friedlander-Mazur} can be reduced to (equivalent to):
\begin{conjecture} \label{FM conj}
\textbf{$\Phi_{1,4}$ is surjective.}
\end{conjecture}

We want to study the relations of this conjecture to  Grothendieck standard conjectures that are fundamental conjectures in the theory of algebraic cycles. We refer to \cite{Kleiman} for the statements of standard conjectures. Recall that the standard conjectures reduce to the standard conjecture of the Lefschetz type.
\

We borrow notations from \cite[Section 2.3]{Beilinson}:
\begin{itemize}
\item
$L(W)$: the Lefschetz type standard conjecture for $W$;
\item
$l(W)$: for every $i > 0$, one can find a  finite correspondence on $W$ that yields an isomorphism $H^{\dim W+ i}(W) \cong H^{\dim W- i}(W)$;
\item
$S(W)$: For any $ j > 0$, one can find a finite correspondence $f_j: W \to Y_j$ with $Y_j$ projective smooth of dimension $j$ such that $f_j^*: H^j(Y_j) \to H^j(W)$ is surjective.
\item
Let $L(n)$ be the assertion that $L(W)$ is true for all $W$ of dimension $\leq n$; same for $l(n), S(n)$.
\end{itemize}
Note that for any smooth projective variety $Y$ with $\dim Y\leq 2$, $L(Y)$ is known. See \cite{Lewis} for example. Therefore $L(2)$ holds.
\

In \cite[Section 2.3]{Beilinson}, Beilinson has proved that $S(n)$ implies $L(W)$ for any smooth projective variety $W$ of dimension $n$ via two steps:
\begin{itemize}
\item
$S(n)$ and $L(n-1)$ implies $l(n)$;
\item
$l(W)$ implies $L(W)$.
\end{itemize}
However the same proof of the first step can be used to prove that $S(X)$, where $X$ is a smooth projective threefold, implies $l(X)$ since $L(2)$ holds. Combining this with his proof of the second step $l(X) \Rightarrow L(X)$, we conclude that $S(X) \Rightarrow L(X)$ for a smooth projective threefold $X$.
\

On the other hand, the Friedlander-Mazur conjecture (\ref{FM conj}) holding for $X$ implies that $S(X)$ via the proof of (i) $\Rightarrow$ (iii) in \cite[Proposition 2.2]{Beilinson}.

In summary, we have the following result.
\begin{theorem}
  The Friedlander-Mazur conjecture for a smooth projective threefold $X$ implies that the Grothendieck standard conjecture holds for $X$.
\end{theorem}

It is already known that Grothendieck standard conjecture holds for a projective threefold  $X$  implies that the Friedlander-Mazur conjecture for $X$ (see \cite[\S 4]{Friedlander2}). Hence  the two conjectures are equivalent in dimension three.

We remark that the Friedlander-Mazur conjecture for all smooth projective varieties is equivalent to that the Grothendieck standard conjecture
of the Lefschetz type  holds.
We emphasis here whether the Friedlander-Mazur conjecture for a given smooth projective variety $X$ is equivalent to the Grothendieck standard conjecture of the Lefschetz type  is still open in dimension four or higher.

\section{Examples}
In most earlier literatures, one verified the Grothendieck standard conjecture first, then use it to  check the Friedlander-Mazur conjecture.
For some  threefolds, it is easier to check the  Friedlander-Mazur conjecture than to check the standard conjectures. This is one of the main purpose in this note.

In \cite{Hu}, it has been proved that the Friedlander-Mazur conjecture holds for a smooth projective threefold $X$ with $h^{2,0}(X) = 0$.
 Hence the standard conjectures also hold for such an $X$. For example,  a smooth projective threefold with representable Chow
 group ${\rm Ch}_0(X)$ of zero cycles satisfying $h^{2,0}(X)=0$. Fano threefolds are examples in this case. A Calabi-Yau threefold $X$ also satisfies $h^{2,0}(X)=0$.
 The last two examples are of threefolds with the Kodaira dimension less than three, the Grothendieck standard conjectures have been proved by
 Tankeev \cite{Tankeev} in a different method.

Surely there exist  lots of threefolds $X$ of general type satisfying $h^{2,0}(X)=0$ but not a complete intersection or a product of a projective curve and surface. Concrete examples of this type can be found in  \cite{Cristina} and references therein. They are new examples of threefolds
for which the Grothendieck standard conjectures of the Lefschetz type hold.
\


\end{document}